# ScNeuGM: Scalable Neural Graph Modeling for Coloring-Based Contention and Interference Management in Wi-Fi 7

### v1

@@ Abstract
Carrier-sense multiple access with collision avoidance in Wi-Fi often leads to contention and interference, thereby increasing packet losses. These challenges have traditionally been modeled as a graph, with stations (STAs) represented as vertices and contention or interference as edges. Graph coloring assigns orthogonal transmission slots to STAs, managing contention and interference, e.g., using the restricted target wake time (RTWT) mechanism introduced in Wi-Fi 7 standards. However, legacy graph models lack flexibility in optimizing these assignments, often failing to minimize slot usage while maintaining reliable transmissions. To address this issue, we propose ScNeuGM, a neural graph modeling (NGM) framework that flexibly trains a neural network (NN) to construct optimal graph models whose coloring corresponds to optimal slot assignments. ScNeuGM is highly scalable to large Wi-Fi networks with massive STA pairs: 1) it utilizes an evolution strategy (ES) to directly optimize the NN parameters based on one network-wise reward signal, avoiding exhaustive edge-wise feedback estimations in all STA pairs; 2) ScNeuGM also leverages a deep hashing function (DHF) to group contending or interfering STA pairs and restricts NGM NN training and inference to pairs within these groups, significantly reducing complexity. Simulations show that the ES-trained NN in ScNeuGM returns near-optimal graphs 4–10 times more often than algorithms requiring edge-wise feedback and reduces 25% slots than legacy graph constructions. Furthermore, the DHF in ScNeuGM reduces the training and the inference time of NGM by 4 and 8 times, respectively, and the online slot assignment time by 3 times in large networks, and up to 30% fewer packet losses in dynamic scenarios due to the timely assignments.
Index Terms:
Interference graph, Wi-Fi, graph modeling.
@@ I Introduction
With ever-increasing demands for wireless connectivity, Wi-Fi 802.11 networks have become a critical component of modern communication infrastructures [1, 2, 3]. The evolution of Wi-Fi technology is now driven by demands for connectivity in emerging Industrial Internet of Things (IIoT) applications, including automated guided vehicles, factory automation, and tactile controls [4, 5, 6]. Typical IIoT networks employ multiple stations (STAs) as sensors to periodically update their status to the access points (APs) via the uplink of Wi-Fi networks [7], as shown in Fig. 1. As delayed or lost sensor statuses can cause inaccurate control decisions and lead to accidents [8], the transmission of these status updates requires stringent quality-of-service (QoS) with low latency and high reliability.
Figure 1: Illustration of a Wi-Fi network for IIoT applications.
@@
Recent Wi-Fi 7 standards introduce the restricted target wake time (RTWT) mechanism [12, 13, 14], which allocates exclusive channel access to STAs during specified time slots (or service periods). This approach improves network performance and STAs’ QoS by eliminating contention and interference between STAs in different slots. When applying RTWT, STAs only need to synchronize slot boundaries and avoid frame transmissions across allocated slot boundaries [15]. Due to its relaxed synchronization requirements, RTWT is an attractive candidate for managing contention and interference across large Wi-Fi IIoT networks. Straightforwardly, each STA can be assigned a separate slot to eliminate all contention and interference in the network. However, such a scheme can cause significant delays in large networks, as each STA must wait for the slots of many other STAs to end. STAs can share a slot if their contention is unlikely to disrupt other STAs’ transmissions, or if the interference among them is negligible. Challenges arise in designing a slot assignment scheme that ensures transmission reliability while minimizing the number of slots needed to separate STAs.
@@ I-A Related Works
Several approaches can model contention and interference and design slot assignment schemes in Wi-Fi 7 IIoT networks.
@@ I-A1 Markov Model Approach
I-A2 Heuristic Graph Model Approach
STA-pairwise contention and interference information can be represented as edges between STAs, forming a graph with STAs as vertices. Slot assignment tasks can then be modeled as graph theory problems [19, 20, 21], such as graph cut [19] or graph coloring [20, 21]. These tasks divide STAs into different slots (either in time or frequency) by disconnecting edges between contending or interfering STA pairs. However, these methods construct graphs—for instance, by assigning edge values between STAs—based on fixed rules predefined according to human intuition about how contention and interference affect STAs’ QoS. As a result, these heuristic graph construction methods lack flexibility and cannot be optimized to meet specific STA QoS requirements.
I-A3 Neural Network Model Approach
@@
In Wi-Fi networks, neural network (NN) models can be flexibly trained to make control decisions based on STAs’ QoS and network data, without requiring explicit contention and interference models [22]. Classic fully connected neural networks (FNNs) have pre-determined input and output dimensions, which limits their scalability to networks of varying sizes [23]. Instead, graph neural networks (GNNs) [24, 25], which are scalable to the graph or the network size, can be applied. GNNs can make decisions by repeatedly aggregating features (network states) of neighboring vertices (or STAs) without being constrained by network size. However, GNNs lose STA-pairwise information during the aggregation of neighboring STAs’ features. This limitation makes them ineffective for tasks where the correlation between each STA pair’s decisions is critical, such as separating STA pairs that are heavily contending or interfering [26, 27]. Additionally, since the minimum number of slots is unknown, it is challenging to design a GNN architecture capable of directly determining each STA’s slot [28, 29].
I-A4 Neural Graph Modeling Approach
@@
NNs can be employed to generate graphs, retaining the graph structure as well as the flexibility to be trained [30, 31]. For example, the authors of [32, 33] use policy gradient (PG) [34] to train NNs that output graphs representing activated devices and transmission links in an integrated sensing and communication system and a cyber-physical power system, respectively, where devices and links are represented as vertices and edges. In Wi-Fi networks, the authors of [35] apply deterministic policy gradient (DPG) [36] to NNs that output a graph model representing STAs as vertices, with contention and interference impacts represented as weighted edges. The graph is cut based on edge weights to partition contending or interfering STAs into a fixed number of slots [35]. However, challenges remain in minimizing slot usage and ensuring reliability in Wi-Fi IIoT networks. Moreover, the above work [35] studies the Wi-Fi networks with limited numbers of devices, e.g., up to tens of devices. Note that the number of possible device pairs grows quadratically when the number of devices increases. How to design the neural graph modeling (NGM) scheme that models the edges in these massive pairs in large Wi-Fi networks requires future research.
I-B Our Contributions
@@
This paper studies the scalable neural graph modeling (ScNeuGM) approach to assign RTWT slots in Wi-Fi 7 networks, e.g., serving periodic IIoT status updates. Our objective is to determine the optimal slot assignments that separate highly contending and interfering STAs, minimizing the number of slots required in each period while ensuring transmission reliability. The network state information used for slot assignments includes path losses from STAs to their nearby APs and the locations of the APs, requiring no additional measurement overheads between STAs. We model the Wi-Fi network as a binary-directed graph, where STAs are represented as vertices and the contention and interference impacts between STAs are represented as binary edges. The coloring scheme of this graph is mapped to the RTWT slot assignment, i.e., STAs are assigned to the same slot if they share the same color, and the chromatic number of the graph is the total number of slots (colors). We formulate the NGM task as training an NGM NN to generate the optimal graph model, i.e., defining the edges connecting STAs, whose coloring scheme minimizes the number of slots while ensuring STAs’ transmission reliability. Two fundamental challenges in the NGM are addressed in our ScNeuGM scheme for large Wi-Fi networks: 1) The absence of explicit STA-pairwise feedback on each generated edge, since the graph’s performance is measured STA-wise (STAs’ reliability) or network-wise (the number of slots); 2) The high computational complexity in the NGM when all STA pairs are processed in training and inference. To avoid acquiring STA-pairwise feedback, an evolution strategy (ES) [37, 38] algorithm is applied to directly adjust the NN parameters without evaluating individual edges. To reduce the complexity, a deep hashing function (DHF) [39] groups likely contending and hidden STA pairs, reducing the processed pairs.
@@
  • To the best of our knowledge, we are the first to formulate the slot minimization problem in the RTWT slot assignment as the NGM task. Unlike GNN models [24, 25], the STA-pairwise information is preserved on the modeled graph edges, indicating whether each STA pair will be separated or not. Moreover, unlike heuristic graph models [19, 20, 21] that construct edges using fixed rules, the proposed scheme flexibly trains NNs to model optimal graphs for specific STA QoS requirements. Further, unlike the previous NGM in Wi-Fi [35] that is only applicable for a fixed number of slots, the proposed one minimizes the number of slots needed to satisfy STA QoS.
  • We apply the ES to optimize the NN to return the graph in the NGM in Wi-Fi networks. Existing works [32, 35] use the PG or the DPG to optimize NN parameters, applying the chain rule to compute the gradient of the graph’s performance w.r.t. the edge values, followed by the gradient of the edge values w.r.t. the NN parameters. Unlike these methods [32, 35], the proposed ES directly adjusts the NN parameters based on the graph’s performance, i.e., STAs’ reliability and the number of slots, without estimating edges’ gradient. The NN trained by the ES is $4$-$10$ times more likely to return near-optimal graphs compared to those trained by the PG or the DPG.
  • Also, we apply the DHF to reduce the NGM complexity by focusing on the training and inference within contending or hidden STA pairs in Wi-Fi networks. To identify contending and hidden pairs, the DHF embeds each STA state into a hash code, where the bits of two STAs’ codes are likely equal if they form a contending or hidden pair. By querying random bits in the STAs’ codes, two STAs are grouped if their bits are likely equal, i.e., they are likely contending or hidden pairs. In large networks with $1000$ STAs, the DHF reduces the training and inference time of the NGM by up to $4$ and $8$ times, respectively, and the online slot assignment time by up to $3$ times.
  • We implement and deploy the ScNeuGM scheme in a system-level, standard-compliant Wi-Fi simulation platform, NS-3 [40], with RTWT enabled. Simulation results show that the graph returned by the ES-trained NN reduces the number of slots by $25\%$ compared to heuristic graphs designed based on human intuition. Further, in scenarios with mobile STAs, the ScNeuGM achieves up to $30\%$ fewer packet losses by applying the DHF to efficiently group STA pairs, enabling timely graph regeneration to adapt to network dynamics compared to the scheme that processes all STA pairs non-selectively.
@@ I-C Notation and Paper Organization
[…]
The rest of this paper is organized as follows. Section II presents the system model. Section III defines the NGM task for contention and interference management and explains the overall concept of the ScNeuGM framework. Sections IV and V present the ES and the DHF in the ScNeuGM, respectively. Finally, Section VI shows the simulations evaluating the proposed methods, and Section VII concludes this work.
@@ II System Model
This section presents the system model of the Wi-Fi network, including the network configurations, the RTWT mechanism, and the network state information.
II-A Wi-Fi Network Configurations
@@
In the network, STAs are assumed to be sensors or monitors continuously collecting data to track critical information in IIoT applications, such as manufacturing systems, and transmitting the collected data samples to the APs via the uplink, as shown in Fig. 1. We consider a Wi-Fi network consisting of $K$ STAs and $A$ APs, as illustrated in Fig. 2. The APs are located in a two-dimensional space, and each AP’s location is known in advance as $\mathrm{loc}_{a}\triangleq(x_{a},y_{a})$, where $x_{a}$ and $y_{a}$ are the coordinates of AP $a$ along the $x$ and $y$ axes, respectively, with $a=1,\dots,A$. All STAs and APs are assumed to operate on the same channel with a bandwidth $B$ in Hertz (Hz). The transmission power of the STAs is $\mathbf{P}_{\mathrm{0}}$ in milliwatts (mW), and the noise power spectral density is $\mathbb{N}_{\mathrm{0}}$ in mW/Hz. The path loss from STA $k$ to AP $a$ is denoted as $s_{k,a}$ in dB, $\forall k,a$. Each STA is associated with the AP that has the minimum path loss to the STA, and the associated AP receives and collects the STA’s transmitted packets. We denote the associated AP of STA $k$ as $\hat{a}_{k}$, i.e.,
$\hat{a}_{k}\triangleq\arg\min_{a}s_{k,a},\ \forall k.$ (1)
We assume that the same data-collection mechanism is used in all STAs over time and that STAs encapsulate each collected data sample into a packet with a constant format, i.e., each STA has the same packet arrival process with a packet size of $L$ bits. Let $d_{k}$ (in seconds) denote the transmission duration of an $L$-bit packet for STA $k$, which depends on the modulation and coding scheme (MCS). Each STA’s MCS is configured to be the highest one for which the decoding error probability without interference, $\epsilon_{k}$, is below a threshold $\epsilon_{\max}=10^{-5}$. The decoding error probability $\epsilon_{k}$ for short packets can be approximated as [41]
@@
where $\phi_{k}=(\mathbf{P}_{\mathrm{0}}/s_{k,\hat{a}_{k}})/(\mathbb{N}_{\mathrm{0}}B)$ is the signal-to-noise-ratio (SNR) of STA $k$ without interference ($s_{k,\hat{a}_{k}}$ here is converted to the decimal scale). In (2), $f_{Q}$ is the tail distribution of the standard normal distribution, and $V$ is the channel dispersion as $V=1-{1}/{\big{[}1+\phi_{k}\big{]}^{2}}$ [41] in (2). Note that we assume that $\mathbf{P}_{\mathrm{0}}$, $\mathbb{N}_{\mathrm{0}}$, $B$, and $L$ in (2) are constants in the network, Thus, for each STA $k$, the packet durations $d_{k}$ only depends on the path loss to STAs’ associated APs, $s_{k,\hat{a}_{k}}$, $\forall k$.

### v2

# Scalable Interference Graph Learning for Low-Latency Wi-Fi Networks using Hashing-based Evolution Strategy

@@ Abstract
Wi-Fi 7 introduces the restricted target wake time (RTWT) mechanism, which is vital for Industrial IoT (IIoT) applications requiring periodic, reliable, and low-latency communication. RTWT enables deterministic channel access by assigning scheduled transmission slots to stations (STAs), minimizing contention and interference. However, determining efficient RTWT slot assignments remains challenging in dense networks, where conventional interference graph-based models lack flexibility and scalability. To overcome this, we propose a scalable interference graph learning (IGL) framework that learns optimal interference graph representations for graph coloring-based RTWT scheduling. The IGL leverages an evolution strategy (ES) to train a neural network (NN) using a single network-wide reward, avoiding costly edge-wise feedback. Furthermore, a deep hashing function (DHF) groups interfering STAs, limiting training and inference to relevant subsets and greatly reducing complexity. Simulation results demonstrate that the proposed IGL improves slot efficiency by up to 25%, reduces packet losses by up to 30% in dynamic environments. Thanks to DHF, it also reduces the training and inference time of IGL by 4 and 8 times, respectively, and the online slot assignment time by 3 times in large networks.
@@ I Introduction
With ever-increasing demands for wireless connectivity, Wi-Fi 802.11 networks have become a vital component of modern communication infrastructure [1, 2, 3]. The evolution of Wi-Fi technology is now driven by demands for connectivity in emerging Industrial Internet of Things (IIoT) applications, including automated guided vehicles, factory automation, and tactile controls [4, 5, 6]. In smart-factory IIoT deployments, hundreds or even thousands of sensors monitor machine health, environmental conditions, and inventory, and must report to a central controller with high timeliness and reliability to support real-time monitoring and control. Typical IIoT networks therefore employ multiple stations (STAs) as sensors that periodically upload status updates to access points (APs) over Wi-Fi uplink channels [7], as shown in Fig. 1. Since delayed or missing status updates can cause inaccurate control decisions and potential accidents [8], these transmissions require stringent quality-of-service (QoS), particularly low latency and high reliability. For example, missing temperature or pressure updates in chemical processes can delay corrective actions and escalate into hazardous events, including explosions [9]; similarly, delayed or lost perception updates in automated guided vehicles can reduce obstacle awareness and increase collision and equipment-damage risks [10].
Figure 1: Illustration of a Wi-Fi network for IIoT applications with the proposed interference graph learning scheme using the hashing-based evolution strategy.
@@
Recent Wi-Fi 7 standards introduce the restricted target wake time (RTWT) mechanism [14, 15, 16], which allocates exclusive channel access to STAs during specified time slots (or service periods). This approach improves network performance and STAs’ QoS by eliminating contention and interference between STAs in different slots. When applying RTWT, STAs only need to synchronize slot boundaries and avoid frame transmissions across allocated slot boundaries [17]. This implies that RTWT has a more relaxed synchronization requirement than orthogonal frequency-division multiple access (OFDMA) [1], which requires symbol-level synchronization among STAs [11]. A straightforward approach is to assign each STA a dedicated slot, which completely eliminates contention and interference. However, this can lead to significant delays in large networks, as each STA must wait for many others to transmit. This issue can be mitigated by allowing slot sharing when STAs are unlikely to interfere with each other’s transmissions. The key challenge, therefore, is to design a slot assignment scheme that ensures reliable communication while minimizing the number of slots used.
@@ I-A Related Works
Several approaches have been proposed to model contention and interference and to design slot assignment schemes for Wi-Fi 7 IIoT networks. We briefly discuss a few key methods.
@@ I-A1 Markov Model Approach
I-A2 Neural Network Model Approach
@@
In Wi-Fi networks, neural network (NN) models can be flexibly trained to make control decisions based on STAs’ QoS and network data, without requiring explicit contention and interference models [21]. Classic fully connected neural networks (FNNs) have pre-determined input and output dimensions, which limits their scalability to networks of varying sizes [22]. Instead, graph neural networks (GNNs) [23, 24], which are scalable to the graph or the network size, can be applied. GNNs can make decisions by repeatedly aggregating features (network states) of neighboring vertices (or STAs) without being constrained by network size. When applied to the separation of contending or interfering STAs in Wi-Fi networks, the aggregation merge the features of neighboring STAs, confusing the differences in interference impacts among individual STA pairs. This makes GNNs ineffective to separate the exact STA pairs that are heavily contending or interfering [25, 26] by assigning orthogonal slots. Additionally, since the minimum number of slots is unknown, it is challenging to design a GNN architecture capable of directly determining each STA’s slot [27, 28].
I-A3 Human Intuition-Defined Interference Graph Approach
STA-pairwise contention and interference information can be represented as edges between STAs, forming a graph with STAs as vertices. Slot assignment tasks can then be modeled as graph theory problems [29, 30, 31], such as graph cut [29] or graph coloring [30, 31]. These tasks divide STAs into different slots (either in time or frequency) by disconnecting edges between contending or interfering STA pairs. However, these methods construct graphs—for instance, by assigning edge values between STAs—based on fixed rules predefined according to human intuition about how contention and interference affect STAs’ QoS. As a result, these graph construction methods lack flexibility and cannot be optimized to meet specific STA QoS requirements.
I-A4 Interference Graph Learning Approach
@@
NNs can be employed to generate graphs, retaining the graph structure as well as the flexibility to be trained [32, 33]. For example, in [34] and [35], the policy gradient (PG) [36] is to train NNs that output graphs representing activated devices and transmission links in an integrated sensing and communication system and a cyber-physical power system, respectively, where devices and links are represented as vertices and edges, respectively. In Wi-Fi networks, the authors of [37] apply deterministic policy gradient (DPG) [38] to NNs that output an interference graph model representing STAs as vertices, with contention and interference impacts represented as weighted edges. The graph is cut based on edge weights to partition contending or interfering STAs into a fixed number of slots [37]. However, challenges remain in minimizing slot usage and ensuring reliability in Wi-Fi IIoT networks. Moreover, the above work [37] studies the Wi-Fi networks with a limited number of devices, e.g., up to tens of devices. Note that the number of possible device pairs grows quadratically when the number of devices increases. Thus, how to design an interference graph learning (IGL) scheme that models the edges in these massive pairs in large Wi-Fi networks requires future research.
Summary of IGL Challenges
In large networks, each STA pair’s edge decision is barely correlated with overall network performance because many other edges are influencing the performance. Therefore, PG and DPG can hardly return a performance-improving gradient direction, since they need to back-propagate each edge decision’s impact on the network performance to the NN parameters. By contrast, evolution strategies (ESs) [39, 40] directly perturb NN parameters and evaluate NN update direction according to the perturbed parameters’ performance. ESs serve as efficient black-box alternatives to the gradient-based method, especially when the credits on how each edge influences the overall performance cannot be explicitly assigned [41]. Nevertheless, how to use ESs in the IGL for large Wi-Fi networks needs further study.
Furthermore, learning interference graphs in large-scale wireless networks is computationally intensive due to the vast STA pairs, resulting in prolonged training and inference time. This delay leads to outdated interference representations that fail to capture dynamic network environments. To mitigate this, the IGL can focus on processing only the pairs of devices that are likely to contend or interfere, setting other edges to null. Given that interference is caused by the geometric proximity of devices, integrating locality-sensitive hashing (LSH) [42] can efficiently identify nearby interfering neighbors, thereby accelerating the IGL process. However, the application of LSH within the IGL remains open.
I-B Contributions
@@
This paper studies the scalable IGL approach to assign RTWT slots in Wi-Fi 7 networks, e.g., serving periodic IIoT status updates. Our objective is to determine the optimal slot assignments that separate highly contending and interfering STAs, minimizing the number of slots required in each period while ensuring transmission reliability. The network state information used for slot assignments includes path losses from STAs to their nearby APs and the locations of the APs, requiring no additional measurement overheads between STAs. We model the Wi-Fi network as a binary-directed graph, where STAs are represented as vertices and the contention and interference impacts between STAs are represented as binary edges. The coloring scheme of this graph is mapped to the RTWT slot assignment, i.e., STAs are assigned to the same slot if they share the same color, and the chromatic number of the graph is the total number of slots (colors). We formulate the IGL task as training an IGL NN to generate the optimal interference graph, i.e., defining the edges connecting STAs, whose coloring scheme minimizes the number of slots while ensuring STAs’ transmission reliability. Two fundamental challenges in the IGL are addressed in our IGL scheme for large Wi-Fi networks: 1) The credit assignment problem [41] on estimating explicit STA-pairwise feedback on a large number of edges, since the graph’s performance is measured STA-wise (STAs’ reliability) or network-wise (the number of slots); 2) The high computational complexity in the IGL when all STA pairs are processed in training and inference. To avoid acquiring STA-pairwise feedback, an ES algorithm [39, 40] is applied to directly adjust the NN parameters without evaluating individual edges. To reduce the complexity, a deep hashing function (DHF) [42] groups likely contending and hidden STA pairs, reducing the processed pairs in the IGL.
@@
  • To the best of our knowledge, we are the first to formulate the slot minimization problem in the RTWT slot assignment as the IGL task. Unlike GNN models [23, 24], the IGL preserves the STA-pairwise interference information on the graph edges, indicating whether each STA pair will be separated or not. Moreover, unlike human intuition-based graph modeling [29, 30, 31] that constructs edges using fixed rules, the proposed scheme flexibly trains NNs to construct optimal graphs for specific STA QoS requirements. Further, unlike the previous IGL in Wi-Fi [37] that is only applicable for a fixed number of slots, the proposed one minimizes the number of slots needed to satisfy STA QoS.
  • We apply the ES to train the NN to construct the optimal interference graph, avoiding the edge-wise credit assignment problem. Existing works [34, 35, 37] use the PG or the DPG to optimize NN parameters via back-propagation, requiring the estimation of the individual edge’s gradient for every edge. Unlike these methods [34, 35, 37], the proposed ES directly correlates each NN parameter’s update direction with the graph’s performance, regardless of the number of edges. The NN trained by the ES generates graphs with 44-$10$ times higher reward signals (indicating fewer slots than human intuition-based graphs while maintaining reliable transmissions) than those by the PG/DPG-trained NNs.
  • Also, we apply the DHF to reduce the IGL complexity by focusing on the training and inference within contending or hidden STA pairs in Wi-Fi networks. To identify contending and hidden pairs, the DHF embeds each STA state into a hash code, where the bits of two STAs’ codes are likely equal if they form a contending or hidden pair. By querying random bits in the STAs’ codes, two STAs are grouped if their bits are likely equal, i.e., they are likely contending or hidden pairs. In large networks with $1000$ STAs, the DHF reduces the training and inference time of the IGL by up to $4$ and $8$ times, respectively, and the online slot assignment time by up to $3$ times.
  • We implement and deploy the IGL scheme in a system-level, standard-compliant Wi-Fi simulation platform, NS-3 [43], with RTWT enabled. Simulation results show that the graph returned by the ES-trained NN reduces the number of time slots by $25\%$ compared to graphs designed based on human intuition at the cost of slightly increased QoS violations, e.g., additional $1\%$ packet loss. Further, in online scenarios with mobile STAs, the proposed IGL achieves up to $30\%$ fewer packet losses by applying the DHF to efficiently group STA pairs, enabling timely graph regeneration to adapt to network dynamics compared to the scheme that processes all STA pairs non-selectively.
@@ I-C Notation and Paper Organization
The $i$-th element of a vector $\mathbf{x}$ is denoted as $x_{i}$. The $j$-th element of the $i$-th row of a matrix $\mathbf{X}$ is denoted as $X_{i,j}$. We define the elements of $\mathbf{X}$ as $\mathbf{X}\triangleq[X_{i,j}\mid X_{i,j}=(\cdot)]$, where $(\cdot)$ represents the expression defining $\mathbf{X}$’s elements. $\operatorname{\bf diag}{\mathbf{X}}$ denotes the diagonal elements of $\mathbf{X}$. $\mathbf{1}_{\{\cdot\}}$ is an indicator that equals $1$ if $\{\cdot\}$ is true and 0 otherwise. $\land$ is logical AND.
The rest of this paper is organized as follows. Section II presents the system model. Section III defines the IGL task for contention and interference management and explains the overall concept of the IGL framework. Sections IV and V present the ES and the DHF in the IGL, respectively. Finally, Section VI shows the simulations evaluating the proposed methods, and Section VII concludes this work.
@@ II System Model
This section presents the system model of the Wi-Fi network, including background on the RTWT mechanism.
II-A Wi-Fi Network with Multi-AP Coordination
@@
In the network, STAs are assumed to be sensors or monitors continuously collecting data to track critical information in IIoT applications, such as manufacturing systems, and transmitting the collected data samples to the APs via the uplink, as shown in Fig. 1. We consider a Wi-Fi network consisting of $K$ STAs and $A$ APs, as illustrated in Fig. 2. The APs are located in a two-dimensional space, and each AP’s location is known in advance as $\mathrm{loc}_{a}\triangleq(x_{a},y_{a})$, where $x_{a}$ and $y_{a}$ are the coordinates of AP $a$ along the $x$ and $y$ axes, respectively, with $a\in\{1,\dots,A\}$. All STAs and APs are assumed to operate on the same channel with a bandwidth of $B$ in Hertz (Hz). The transmission power of the STAs is $\mathbf{P}_{\mathrm{0}}$ in milliwatts (mW), and the noise power spectral density is $\mathbb{N}_{\mathrm{0}}$ in mW/Hz.
The APs are coordinated by a single controller that manages the STAs’ connections. The controller runs at an edge server that has real-time communication with all APs, via which it receives network measurements and sends control decisions from and to APs. The coordinated APs form an extended service set [11] where STAs can roam between APs without resending re-association requests to APs. Each STA is associated with the AP that has the minimum path loss to the STA, and the associated AP receives and collects the STA’s transmitted packets. We denote the path loss from STA $k$ to AP $a$ as $s_{k,a}$ in dB, $\forall k,a$ and the associated AP of STA $k$ as $\hat{a}_{k}$, i.e.,
$\hat{a}_{k}\triangleq\operatorname*{\arg\!\min}_{a\in\{1,\ldots,A\}}s_{k,a}.$ (1)
Each AP keeps measuring its path losses to the STAs based on the STAs’ packets detected and received at the AP. The measured path losses will be updated to the controller that decides which AP should be associated with the STA and process the STA’s packets. We assume that the same data-collection mechanism is used in all STAs over time and that STAs encapsulate the collected data sample into a packet with a constant format, i.e., each STA has the same packet arrival process with a packet size of $L$ bits.
The controller will notify each STA of the AP that it is associated with by instructing the APs to send a notification message. The message contains the path loss from the STA to the associated AP whenever the association changes or the path loss changes significantly. Then, each STA can adjust the modulation and coding scheme (MCS) of its packet transmissions according to the new path loss. Specifically, let $d_{k}$ (in seconds) denote the transmission duration of an $L$-bit packet for STA $k$, which depends on the MCS. Each STA’s MCS is configured to be the highest one for which the decoding error probability without interference, $\epsilon_{k}$, is below a threshold, say $\epsilon_{\max}=10^{-5}$. The decoding error probability $\epsilon_{k}$ for short packets can be approximated as [44]
@@
where $\phi_{k}=\frac{1}{s_{k,\hat{a}_{k}}}\frac{\mathbf{P}_{\mathrm{0}}}{\mathbb{N}_{\mathrm{0}}B}$ is the signal-to-noise-ratio (SNR) at STA $k$ without interference ($s_{k,\hat{a}_{k}}$ here is converted to the decimal scale). In (2), $f_{Q}$ is the tail distribution of the standard normal distribution, and $V$ is the channel dispersion as $V=1-\frac{1}{(1+\phi_{k})^{2}}$ [44] in (2). Note that we assume that $\mathbf{P}_{\mathrm{0}}$, $\mathbb{N}_{\mathrm{0}}$, $B$, and $L$ in (2) are constants in the network. Thus, for each STA $k$, the packet duration $d_{k}$ only depends on the path loss to STAs’ associated APs, $s_{k,\hat{a}_{k}}$, $\forall k$. Throughout the paper, we assume that $d_{k}$ is bounded and less than the slot duration.